# Convergence Analysis of Non-Strongly-Monotone Stochastic Quasi-Variational Inequalities

### v3

@@ 2.3 Saddle Point Problems with Coupling Constraints
Recently, minimax optimization problems, i.e., $\min_{u\in U}\max_{w\in W}f(u,w)$, have received considerable attention due to their relevance to a wide range of machine learning (ML) problems such as reinforcement learning, generative adversarial networks (GAN), fairness in machine learning, and generative adversarial imitation learning. Convex-concave minimax problems can be also viewed from a game theory perspective in which the objective function severs as a payoff function and one player aims to minimize the payoff while the second player’s goal is to maximize it. In this context, a saddle point solution $(u^{*},w^{*})$ serves as both a minimum of $f$ in the $u$-direction and a maximum of $f$ in the $w$-direction. Here, $w^{*}$ represents the inner player’s best response to their opponent’s strategy $u^{*}$, and a saddle point $(u^{*},w^{*})$ is also called a Nash equilibrium (NE).
@@ 4.1 Notations
4.2 Assumptions and Technical Lemmas
@@
Next, we state our main assumptions considered in the paper.
@@
There exists $\nu>0$ such that $\mathbb{E}[\bar{w}_{k,N_{k}}\break\mid\mathcal{F}_{k}]=0$ and $\mathbb{E}[\|\bar{w}_{k,N_{k}}\|^{2}\mid\mathcal{F}_{k}]\leq\tfrac{\nu^{2}}{N_{k}}$ holds almost surely for all $k$, where $\mathcal{F}_{k}\triangleq\sigma\{x_{0},x_{1},\ldots,x_{k-1}\}$ and $\bar{w}_{k,N_{k}}\triangleq\tfrac{1}{N_{k}}{\sum_{j=1}^{N_{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(G(x_{k},\xi_{j,k})-F(x_{k}))}}$.
@@ 5 Proposed Method
There exists $\gamma>0$ such that $\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(x)}[u]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(y)}[u]\|\leq\gamma\|x-y\|$ for all $x,y,u\in X$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(x)}[u]={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{m(x)+K}[u]=m(x)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(x)],\quad\forall x,u\in X.$
@@
$\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(x)}[u]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K(y)}[u]\|$ $=\|m(x)-m(y)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(x)]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(y)]\|$
$\leq\|m(x)-m(y)\|+\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(x)]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}}_{K}[u-m(y)]\|$
$\leq 2\|m(x)-m(y)\|\leq 2\tilde{\gamma}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|x-y\|}.$
@@
Input: $x_{0}\in X$, $\eta>0$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T>0}$, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{b_{k}\}_{k}}$, $\{\alpha_{k}\}_{k}$ and Algorithm $\mathcal{M}$ satisfying Assumption 4;
@@ 5.1 Convergence Analysis
±$\quad\pm\alpha_{k}\mathcal{P}_{K(x_{k})}\left[u_{k}-\eta(F(u_{k})+\bar{w}_{k,N_{k}})\right]\|$
@@
where we assume that the product is 1 when there are no terms in the multiplication, i.e., $\prod_{j=i}^{k-1}(1-q_{j+1})=1$ if $i>k-1$.
@@
Since $\rho\geq 1-q$ and $q\in(0,1)$, one can easily confirm that $-\frac{(1-q)^{T}}{\rho+q-1}<0$, hence the following holds.
@@
By applying the Cauchy-Schwarz inequality, using the fact that $D\triangleq\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k=0}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\infty}}\break\tfrac{1}{(k+1)\log^{2}(k+2)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq 3.39}$ and rearranging the terms, the desired result is obtained:
@@
(ii) To compute an $\epsilon$-solution, i.e., $\mathbb{E}[\|x_{T}-\bar{x}_{T}\|]\leq\epsilon$, it follows from (9) that $T=\log_{1/\rho}(\bar{D}/\epsilon)$ iterations is required, where $\bar{D}=\|x_{0}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{0}}\|\break+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\eta(\tfrac{\beta\bar{b}\nu^{\prime}+\nu}{\rho+q-1})+\tfrac{\bar{\alpha}D(\beta\bar{b}C^{\prime}+C)}{\rho}}$. Moreover, in Algorithm 1, each iteration requires taking $t_{k}=\tfrac{({k+1})\log^{2}(k+2)}{\rho^{k}}$ inner steps of Algorithm $\mathcal{M}$. Therefore, the total number of inner iterations is
@@ 5.2 Gradient Approach
Input: $x_{0}\in X$, $\eta>0$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T>0}$, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, $\{\alpha_{k}\}_{k}$ and Algorithm ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}$ satisfying Assumption 4;
@@
Note that to obtain convergence rate guarantees, we required the condition $\gamma+\sqrt{1-\frac{\mu_{F}^{2}}{L^{2}}}<1$, which is problem-dependent and commonly appears in prior works on deterministic QVI problems with convergence guarantees (e.g., see Theorem 2 and 3 in [31] and Theorem 5 in [35]). Whether convergence rate guarantees—such as sublinear or linear rates—can be established without relying on this condition remains an open question, particularly due to the added complexity introduced by feasible sets that depend on the decision variable. In the next section, we show that almost sure convergence can be achieved without relying on this structural condition.
5.3 Almost Sure Convergance
@@
[…]
@@
Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by Algorithm 2, and suppose Assumption 1 and 2 hold and $\{N_{k}\}_{k}$ is an increasing sequence, such that $\sum_{k=0}^{\infty}\tfrac{1}{\sqrt{N_{k}}}<\infty$. Then the sequence $\{x_{k}\}$ has at least one limit point, and every limit point is a solution of the SQVI almost surely.

### v4

@@ 2.3 Saddle Point Problems with Coupling Constraints
Recently, minimax optimization problem, i.e., $\min_{u\in U}\max_{w\in W}f(u,w)$, has received considerable attention due to its relevance to a wide range of machine learning (ML) problems such as reinforcement learning, generative adversarial networks (GAN), fairness in machine learning, and generative adversarial imitation learning. Convex-concave minimax problems can be also viewed from a game theory perspective in which the objective function severs as a payoff function and one player aims to minimize the payoff while the second player’s goal is to maximize it. In this context, a saddle point solution $(u^{*},w^{*})$ serves as both a minimum of $f$ in the $u$-direction and a maximum of $f$ in the $w$-direction. Here, $w^{*}$ represents the inner player’s best response to their opponent’s strategy $u^{*}$, and a saddle point $(u^{*},w^{*})$ is also called a Nash equilibrium (NE).
@@ 4.1 Notations
4.2 Quadratic Growth Property
@@
Next, we consider a class of generalized Nash equilibrium problems that can be reformulated as QVIs. For this class, we explicitly characterize the equilibrium set $X^{*}$ and show that the associated QVI operator satisfies the QG property.
Class of QVIs with Quadratic Growth. Consider a generalized Nash game in which each player $i\in\{1,\cdots,N\}$ solves the following optimization problem:
$\min_{x_{i}\in X_{i}}f_{i}(A_{i}x_{i})+c_{i}^{T}x^{*}_{(-i)}$
$\mbox{s.t.}\quad B_{i}x_{i}+C_{i}x_{(-i)}^{*}=d_{i},$
where $X_{i}\triangleq\{x_{i}\mid D_{i}x_{i}\leq q_{i}\}$ is a polyhedral set, function $f_{i}$ is strongly convex, $x^{*}_{(-i)}$ denotes the equilibrium decisions of the other players. Note that the problem may not be strongly monotone, when $A$ is not full-rank. This game can be equivalently reformulated as a quasi variational inequality (QVI) problem.
Let $x=(x_{1},\ldots,x_{N})$ denote the joint decision vector, and define the mapping
$F(x)=\big(A_{1}^{\top}\nabla f_{1}(A_{1}x_{1}),\ldots,A_{N}^{\top}\nabla f_{N}(A_{N}x_{N})\big).$
Each player’s feasible set depends on the others’ strategies through the shared constraints $(B_{i}x_{i}+C_{i}x_{(-i)}=d_{i})$, leading to the coupling feasible region
$K_{i}(x_{(-i)})=\{x_{i}\in X_{i}:B_{i}x_{i}+C_{i}x_{(-i)}=d_{i},\ \forall i=\{1,\ldots,N\}\}.$
The Nash equilibrium conditions then correspond to finding $x^{*}\in K(x^{*})=\Pi_{i}K_{i}(x_{(-i)}^{*})$ such that $\langle F(x^{*}),x-x^{*}\rangle\geq 0,\quad\forall x\in K(x^{*}),$ which is a QVI problem. Inspired by Necoara, Nesterov, and Glineur [2], we show that the QVI reformulation satisfies QG property if there exits $\eta_{i}>0$ such that $B_{i}^{T}B_{i}\preceq\eta_{i}A_{i}^{T}A_{i}$ for any $i$.
First we show that for each $i$ there exists a unique pair ($t_{i}^{*},T_{i}^{*})$ such that $A_{i}x_{i}^{*}=t_{i}^{*}$ and $A_{i}^{\top}\nabla f_{i}(A_{i}x_{i}^{*})=T_{i}^{*}$. Let $x_{i}^{*1},x_{i}^{*2}$ be two optimal strategy of player $i$, and define $\phi_{i}(x_{i},x_{(-i)})\triangleq f_{i}(A_{i}x_{i})+c_{i}^{T}x^{*}_{(-i)}$, then using convexity of $\phi_{i}(\cdot,x_{(-i)})$ and the fact that $x_{i}^{*}$ is optimal, i.e., $\phi_{i}(x_{i}^{*},x^{*}_{(-i)})\leq\phi_{i}(x_{i},x^{*}_{(-i)})$ for all $x_{i}\in K_{i}(x^{*}_{(-i)})$. It is also clear that $\tfrac{x_{i}^{*1}+x_{i}^{*2}}{2}$ is a feasible point, hence, we have that $\phi_{i}(\tfrac{x_{i}^{*1}+x_{i}^{*2}}{2},x^{*}_{(-i)})=\tfrac{\phi_{i}(x_{i}^{*1},x^{*}_{(-i)})}{2}+\tfrac{\phi_{i}(x_{i}^{*2},x^{*}_{(-i)})}{2}$, which implies that $f_{i}(\tfrac{A_{i}x_{i}^{*1}+A_{i}x_{i}^{*2}}{2})=\tfrac{f_{i}(A_{i}x_{i}^{*1})+f_{i}(A_{i}x_{i}^{*2})}{2}$. On the other hand, since $f_{i}$ is strongly monotone with modulus $\mu_{i}$, we have that
$f_{i}(\tfrac{A_{i}x_{i}^{*1}+A_{i}x_{i}^{*2}}{2})\leq\tfrac{f_{i}(A_{i}x_{i}^{*1})}{2}+\tfrac{f_{i}(A-ix_{i}^{*2})}{2}-\tfrac{\mu_{i}}{8}\|A_{i}x_{i}^{*1}-A_{i}x_{i}^{*2}\|^{2}.$
Hence, we conclude that $A_{i}x_{i}^{*1}=A_{i}x_{i}^{*2}$.
Therefore, for each player $i$, the optimal solution set can be charactrized as $X_{i}^{*}(x^{*}_{(-i)})=\{x_{i}\mid A_{i}x_{i}=t_{i}^{*},B_{i}x_{i}=d_{i}-C_{i}x_{(-i)}^{*},D_{i}x_{i}\leq q_{i}\}$. Since $X_{i}^{*}(x^{*}_{(-i)})$ is a polyhedron and is nonempty, then from Hoffman inequality we have that there exists some positive constant depending on the matrices $A_{i}$, $B_{i}$ and $D_{i}$ describing the polyhedral set $X_{i}^{*}(x^{*}_{(-i)})$, i.e., $\theta_{i}(A_{i},B_{i},D_{i})>0$, such that for any $x_{i}\in X_{i}$:
$\|x_{i}-\bar{x}_{i}\|^{2}$ $\leq\theta_{i}^{2}(A_{i},B_{i},D_{i})\left(\|A_{i}x_{i}-A_{i}\bar{x}_{i}\|^{2}+\|B_{i}x_{i}-B_{i}\bar{x}_{i}\|^{2}\right)$
$\leq\eta_{i}\theta_{i}^{2}(A_{i},B_{i},D_{i})\|\|A_{i}x_{i}-A_{i}\bar{x}_{i}\|^{2},$
where $\bar{x}_{i}=\mathcal{P}_{X_{i}^{*}(x^{*}_{(-i)})}(x_{i})$ and we used the assumption that $B_{i}^{T}B_{i}\preceq\eta_{i}A_{i}^{T}A_{i}$. Using this result, one can show that gradient of each player objective satisfies quadratic growth property:
$\langle A_{i}^{T}\nabla f_{i}(A_{i}x_{i})-A_{i}^{T}\nabla f_{i}(A_{i}\bar{x}_{i}),x_{i}-\bar{x}_{i}\rangle$ $=\langle\nabla f_{i}(A_{i}x_{i})-\nabla f_{i}(A_{i}\bar{x}_{i}),A_{i}x_{i}-A_{i}\bar{x}_{i}\rangle$
$\geq\mu_{i}\|A_{i}x_{i}-A_{i}\bar{x}_{i}\|$
$\geq\tfrac{\mu_{i}}{\sqrt{\eta_{i}}\theta_{i}(A_{i},B_{i},D_{i})}\|x_{i}-\bar{x}_{i}\|.$
Therefore, summing across $i$, the full operator $F$ satisfies: $\langle F(x)-F(\bar{x}),x-\bar{x}\rangle\geq\mu\|x-\bar{x}\|^{2}$ where $\mu=\min_{i}\{\tfrac{\mu_{i}}{\sqrt{\eta_{i}}\theta_{i}(A_{i},B_{i},D_{i})}\}$ and $\bar{x}\in X^{*}=\Pi_{i=1}^{N}X_{i}^{*}(x^{*}_{(-i)})$.
4.3 Assumptions and Technical Lemmas
Now, we state our main assumptions considered in the paper.
@@ 4.3 Assumptions and Technical Lemmas
There exists $\nu>0$ such that $\mathbb{E}[\bar{w}_{k,N_{k}}\mid\mathcal{F}_{k}]=0$ and $\mathbb{E}[\|\bar{w}_{k,N_{k}}\|^{2}\mid\mathcal{F}_{k}]\leq\tfrac{\nu^{2}}{N_{k}}$ holds almost surely for all $k$, where $\mathcal{F}_{k}\triangleq\sigma\{x_{0},x_{1},\ldots,x_{k-1}\}$ and $\bar{w}_{k,N_{k}}\triangleq\tfrac{1}{N_{k}}{\sum_{j=1}^{N_{k}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(G(x_{k},\xi_{j,k})-F(x_{k}))}}$.
@@ 5 Proposed Method
There exists $\gamma>0$ such that $\|{\mathcal{P}}_{K(x)}[u]-{\mathcal{P}}_{K(y)}[u]\|\leq\gamma\|x-y\|$ for all $x,y,u\in X$.
@@
${\mathcal{P}}_{K(x)}[u]={\mathcal{P}}_{m(x)+K}[u]=m(x)+{\mathcal{P}}_{K}[u-m(x)],\quad\forall x,u\in X.$
@@
$\|{\mathcal{P}}_{K(x)}[u]-{\mathcal{P}}_{K(y)}[u]\|$ $=\|m(x)-m(y)+{\mathcal{P}}_{K}[u-m(x)]-{\mathcal{P}}_{K}[u-m(y)]\|$
$\leq\|m(x)-m(y)\|+\|{\mathcal{P}}_{K}[u-m(x)]-{\mathcal{P}}_{K}[u-m(y)]\|$
$\leq 2\|m(x)-m(y)\|\leq 2\tilde{\gamma}{\|x-y\|}.$
@@
Input: $x_{0}\in X$, $\eta>0$, T>0{T>0}, $\{N_{k}\}_{k}$, $\{t_{k}\}_{k}$, $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{b_{k}\}_{k}$}, $\{\alpha_{k}\}_{k}$ and Algorithm $\mathcal{M}$ satisfying Assumption 4;
@@ 5.1 Convergence Analysis
$\quad\pm\alpha_{k}\mathcal{P}_{K(x_{k})}\left[u_{k}-\eta(F(u_{k})+\bar{w}_{k,N_{k}})\right]\|$
@@
where we assume that the product is 1 when there are no terms in the multiplication, i.e., $\prod_{j=i}^{k-1}(1-q_{j+1})=1$ if i>k1i>k-1.
@@
Since $\rho\geq 1-q$ and $q\in(0,1)$, one can easily confirm that $\frac{(1-q)^{T}}{\rho+q-1}<0$, hence the following holds.
@@
By applying the Cauchy-Schwarz inequality, using the fact that $D\triangleq\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k=0}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\infty}}\penalty-10000\tfrac{1}{(k+1)\log^{2}(k+2)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq 3.39}$ and rearranging the terms, the desired result is obtained:
@@
(ii) To compute an $\epsilon$-solution, i.e., $\mathbb{E}[\|x_{T}-\bar{x}_{T}\|]\leq\epsilon$, it follows from (9) that $T=\log_{1/\rho}(\bar{D}/\epsilon)$ iterations is required, where $\bar{D}=\|x_{0}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{x}_{0}}\|\penalty-10000+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\alpha}\eta(\tfrac{\beta\bar{b}\nu^{\prime}+\nu}{\rho+q-1})+\tfrac{\bar{\alpha}D(\beta\bar{b}C^{\prime}+C)}{\rho}}$. Moreover, in Algorithm 1, each iteration requires taking $t_{k}=\tfrac{({k+1})\log^{2}(k+2)}{\rho^{k}}$ inner steps of Algorithm $\mathcal{M}$. Therefore, the total number of inner iterations is
@@ 5.2 Gradient Approach
[…]
@@
Note that to obtain convergence rate guarantees, we required the condition $\gamma+\sqrt{1-\frac{\mu_{F}^{2}}{L^{2}}}<1$, which is problem-dependent and commonly appears in prior works on deterministic QVI problems with convergence guarantees (e.g., see Theorem 2 and 3 in [30] and Theorem 5 in [34]). Whether convergence rate guarantees—such as sublinear or linear rates—can be established without relying on this condition remains an open question, particularly due to the added complexity introduced by feasible sets that depend on the decision variable. In the next section, we show that almost sure convergence can be achieved under a weaker requirement, which becomes less restrictive and easier to satisfy.
5.3 Almost Sure Convergence
@@
where $\mathbb{E}\left[\omega_{t+1}\mid\mathcal{F}_{t}\right]$ represents the conditional expectation for the given $\{\omega_{0}...,\omega_{t}\}$, $\{\eta_{0},...,\eta_{t}\},$ $\{\upsilon_{0},...,\upsilon_{t}\}$, $\{\psi_{0},...,\psi_{t}\}$, and $\sum_{t=0}^{\infty}\eta_{t}<\infty$, $\sum_{t=0}^{\infty}\upsilon_{t}<\infty$ , then we have:  $\text{ (i) }\omega_{t}\rightarrow\omega$ almost surely, and  $\text{ (ii) }\sum_{t=0}^{\infty}\psi_{t}<\infty$.
@@
Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by Algorithm 2, and suppose Assumption 1 and 2 hold and $\{N_{k}\}_{k}$ is an increasing sequence, such that $\sum_{k=0}^{\infty}\tfrac{1}{{N_{k}}}<\infty$. Choose $\eta=\frac{1.8\mu_{F}}{L^{2}}$, and $\gamma\leq 45(\sqrt{1+0.04\frac{\mu_{F}^{2}}{L^{2}}}-1)$, and $\alpha_{k}=\tfrac{0.01}{(k+1)^{0.5+\delta}}$, for some $\delta>0$.Then the sequence $\{x_{k}\}$ has at least one limit point, and every limit point is a solution of the SQVI almost surely.
Note that the upper-bound for $\gamma$ in Theorem 3 is greater than the upper-bound in Theorem 2, i.e., $45(\sqrt{1+0.04\frac{\mu_{F}^{2}}{L^{2}}}-1)>1-\sqrt{1-\frac{\mu_{F}^{2}}{L^{2}}}$, when $\frac{\mu_{F}}{L}\leq 0.992$, which is the case for most real-world problems, as $\frac{\mu_{F}}{L}$ is typically much smaller than 11.